# SEE-2-SOUND  [Uncaptioned image] : Zero-Shot Spatial Environment-to-Spatial Sound

### v1

@@ 4.2 Audio Generation
We are now interested in learning from the set of image patches $\mathbf{I}_{\mathbf{p}}$ and an optional text prompt $\mathbf{c}$, a set of 4D representations, $(x,y,z,\mathbf{A}^{T\times F})$ representing the positions and audio. We generate mono audio clips conditioned on each of the image patches in ${\mathcal{R}}$ and the text conditioning $\mathbf{c}$ using a pre-trained CoDi model [72]. We thus have the 4D representation $(x,y,z,\mathbf{A}^{T\times F})$ for each of the individual image patches in $\{{\mathcal{R}}_{1},{\mathcal{R}}_{2},{\mathcal{R}}_{3},\ldots\}$.
4.3 Generating 5.1 Surround Sound Spatial Audio
@@
Figure 3: Overview of the simulation. We show a schematic of the virtual “ShoeBox” room utilized in our 5.1 surround sound simulation. The diagram presents the three-dimensional arrangement of multiple audio sources (labeled Source 1, Source 2, and Source 3) and the standardized microphone placement for the channels Front-Left (FL), Front-Right (FR), Center (C), Surround-Left (SL), Surround-Right (SR), and Low-Frequency Effects (LFE).
@@
The goal of surround sound spatial audio generation is to generate ambisonic channels $\phi_{N}(t)$ from different non-spatial audios and positions $(x,y,z,\mathbf{A}^{T\times F})$. We first construct a virtual acoustic environment, specifically a “ShoeBox” room [62]. We place the listener centrally at the ground level of this room, serving as a reference point for audio projection and perception. For the source and microphone placements, we adhere to the ITU-R BS 775 configuration [3], which is essential for aligning our simulations with the standard 5.1 Surround Sound setups. This configuration includes designated positions for the Front-Left (FL), Front-Right (FR), Center (C), Surround-Left (SL), Surround-Right (SR), and Low-Frequency Effects (LFE) channels [65].
A critical component of our simulation involves computing the Room Impulse Responses (RIRs) using the Image Source Method (ISM) [3]. We now calculate the impulse response $h(t)$ by accounting for contributions from multiple virtual sources that arise from reflections off the room’s boundaries. For a source at position $\mathbf{s}$ and a receiver at position $\mathbf{r}$, the impulse response is expressed as
$h(t)=\sum_{m}\alpha_{m}\delta(t-\tau_{m}),$ (4)
where $\tau_{m}=\frac{\|\mathbf{r}-\mathbf{s}_{m}\|}{c}$ denotes the time delay; $c$ represents the speed of sound, and $\alpha_{m}$ are the attenuation factors due to reflections. This model effectively captures both the early reflections and the reverberations, which are crucial for rendering a realistic acoustic environment.
After computing the RIRs, we convolve these responses with the original mono audio tracks to synthesize the spatial audio for each of the 5.1 channels. The convolution integrates the spatial effects into the audio, effectively distributing the sound across the designated channels. The synthesized signal for each channel $ch$ is expressed as
$\phi_{ch}(t)=\mathbf{A}^{T\times F}\ast h_{ch}(t),$ (5)
@@
where $\mathbf{A}^{T\times F}$ denotes the original audio signal in the time-frequency domain. Our simulation also incorporates the interference among audio sources which is a significant factor that differentiates the outcome from scenarios where fewer sources are used and then combined. Thus, the resulting sound field in a full simulation exhibits unique characteristics that cannot be replicated by simply summing the outputs of simulations with fewer sources. This interference results in spatial audio properties that are intrinsically linked to the arrangement $(x,y,z)$ and the number of sources in the complete setup.
5 Experiments
Table 1: Quantitative Evaluation Results. We report average similarity scores between different audios. The similarity metrics are calculated across the method and AviTAR [14] enhanced audios of the methods. Lower MFCC-DTW scores and higher ZCR, Chroma, and Spectral scores indicate higher similarity.
(+ AViTAR [14]) MFCC-DTW $\downarrow$ ZCR $\uparrow$ Chroma $\uparrow$ Spect $\uparrow$
CoDi [72] (baseline) 0.8 ×$\times 10^{-3}$ 0.80 0.70 0.85
SEE-2-SOUND (ours) 0.03 ×$\times 10^{-3}$ 0.95 0.77 0.95
@@ 5 Experiments
To the best of our knowledge, no previous techniques address the problem of spatial sound generation, thus we propose a novel quantitative evaluation technique based on marginal scene guidance (Section 5.1) alongside human evaluation (Section 5.2).
5.1 Quantitative Evaluation
@@
We conduct a quantitative evaluation of the scene guidance of generations produced by our model using various audio similarity metrics. To construct this setup, we first sample images from Laion400M [63] and the web. We then generate audio from these images using our method as well as CoDi [72] which serves as a baseline. We then use the audio outputs and images to generate modified scene-guided audio generated using AViTAR [14]. The use of an unconventional model-based evaluation stems from the lack of existing evaluations or any strong baselines. We report the average audio similarity scores between multiple such generated audio in Table 1. In particular, SEE-2-SOUNDconsistently generates spatial audio that matches the scene’s content as shown by AViTAR [14] only slightly altering the generated audios.
5.2 Human Evaluation
Due to the absence of any traditional metrics, we also use human evaluation to evaluate our approach. To promote future research in this field, we will open-source the evaluation protocol for human evolution. We conduct human evaluations by surveying $43$ human evaluators, for which we use randomly selected images and videos from the internet that are non-iconic in nature. We use this qualitative evaluation technique specifically due to the absence of any baseline methods for spatial generation. Additionally, the task of spatial audio generation is challenging since there are multiple possible ground truth spatial audios for a given image or video. This makes it difficult for us to use any kind of similarity metric. Existing similarity metrics like LPIPS [87] or CLIP score [52] and their audio equivalents that make use of embeddings are also not feasible since we do not train embeddings directly for the spatial generation problem.
Table 2: Human Evaluation Results. We evaluate our model performance across metrics identified through a user study. The hypothesis of our test was that the responses follow a uniform distribution. For the quality questions, we considered 3, 4, and 5 to be correct on a scale of 1-5. For the distance and localization problems, we used the patches and coordinates generated by our approach as correct answers. For adversarial identification and matching problems, we knew the correct answers by the nature of the problem. The significance of human preference is evaluated with $\chi^{2}$-tests at $p<0.05$.
Metric $p$-value ($\downarrow$)
Spatial Audio Quality
@@ 5.2 Human Evaluation
   Overall Localization $4.523\times 10^{-4}$
   Audio Direction Identification $2.880\times 10^{-7}$
   Distance Identification $3.103\times 10^{-2}$
@@
Spatial Audio Quality.
@@
Adversarial Identification.
Evaluators are given a random mix of generated spatial audios and mono audios. They are tasked with discriminating whether the given audio is indeed spatial audio, their accuracy is measured.
Audio Quality.
@@
More details on the tasks in our user study can be found in Appendix B. We evaluate the results for each of these individual sub-tasks and questions and conduct $\chi^{2}$-tests to evaluate significance at $p<0.05$ level.
Furthermore, we demonstrate our outputs in the supplementary from which one can observe that SEE-2-SOUND produces compelling spatial audio that aligns well with videos and images which are either generated from video generation models or taken from the internet. We perform the user testing on $43$ human evaluators and report the metrics for our human evaluation in Table 2. Users indicated a statistically significant preference towards SEE-2-SOUND in terms of spatial audio quality, directions of sound, and audio quality.
6 Discussion
@@
Lastly, inherent in the capabilities of generative models such as SEE-2-SOUND– which excel in producing synchronized and realistic outputs – is a significant concern regarding the potential generation and dissemination of deepfakes. The exploitation of this technology by malicious entities poses a notable risk even though we do not generate speech, as they could fabricate realistic counterfeit audio clips. We condemn any such applications and hope to be able to work on safeguards for such approaches in the future.
7 Conclusion
@@ Acknowledgements
We thank Bhavya Bhatt of the University of Toronto for his help on surveying the existing literature. This research was enabled in part by support provided by the Digital Research Alliance of Canada  [^1] https://alliancecan.ca/. This research was supported in part with Cloud TPUs from Google’s TPU Research Cloud (TRC) [^2] https://sites.research.google/trc/about/. The resources used to prepare this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute  [^3] https://vectorinstitute.ai/partnerships/current-partners/.
@@ Appendix A Details of our Method
We provide the details of our approach with which we demonstrate all of our experiments and results. Our experiments ran on a single A100-80 GPU. Running our approach on an image takes $\sim 12-15$ seconds (non-initial runs; initial runs can take up to $\sim 3$ minutes) for images and $\sim 5$ minutes on videos that are under $20$ seconds like the ones we experiment with. Our method can be easily swapped for other off-the-shelf models however we limit our experiments to these settings. Our human evaluation experiments were also performed without any text prompts to solely determine our framework’s performance.
@@
For source estimation we use a Segment Anything [33] model. We use the ViT-H/16 [20] with $14\times 14$ windowed attention as our image encoder. Following standard approaches we rescale any inputs to $1024\times 1024$ by rescaling and padding. For the decoder, we use the same design and weights as used by Segment Anything [33]. For monocular depth estimation, we use a Depth Anything [84] model. We use ViT-L/14 [20] as our image encoder.
@@
For audio generation we use a CoDi [72] model and use the Video LDM, Audio LDM, and Text LDM components. For inference, we sample the model through $500$ diffusion steps.
@@
We use a sampling rate of $16000$ Hz for audio processing. For simulating the acoustic environment, we utilize a “ShoeBox" room model from Pyroomacoustics [62]. The room dimensions are dynamically calculated based on the image size. Specifically, the dimensions width $w=$ image_width, height $h=$ image_height, and length $l=0.5\cdot$ image_width. In our virtual room, the microphones are placed at the following coordinates for each channel,
@@
Appendix B User Study Prootocol
@@
B.1 Consent Form
Your decision to participate is voluntary and you are free to withdraw at any time without any consequences. You are not required to answer any questions you do not feel comfortable answering.
@@ B.2 Part 1: Familiarization
B.3 Part 2: Rating Spatial Audio with Images
@@
B.4 Part 3: Direction and Distance Identification
@@
B.5 Part 4: Real vs. Fake Identification
You will be shown 2 clips, one of them has spatial audio generated by and another one has mono audio of the same clips. You will be asked to identify which of them is "real" spatial audio and "fake" spatial audio.
@@
B.6 Part 5: Matching Images to Spatial Audio

### v2

@@ 4.2. Audio Generation
We are now interested in learning from the set of image patches $\mathbf{I}_{\mathbf{p}}$ and an optional text prompt $\mathbf{c}$, a set of 4D representations, $(x,y,z,\mathbf{A}^{T\times F})$ representing the positions and audio. We generate mono audio clips conditioned on each of the image patches in ${\mathcal{R}}$ and the text conditioning $\mathbf{c}$ using a pre-trained CoDi model (Tang et al., 2024). We thus have the 4D representation $(x,y,z,\mathbf{A}^{T\times F})$ for each of the individual image patches in $\{{\mathcal{R}}_{1},{\mathcal{R}}_{2},{\mathcal{R}}_{3},\ldots\}$.
4.3. Generating 5.1 Surround Sound Spatial Audio
@@
Figure 3. Overview of the simulation. We show a schematic of the virtual “ShoeBox” room utilized in our 5.1 surround sound simulation. The diagram presents the three-dimensional arrangement of multiple audio sources (labeled Source 1, Source 2, and Source 3) and the standardized microphone placement for the channels Front-Left (FL), Front-Right (FR), Center (C), Surround-Left (SL), Surround-Right (SR), and Low-Frequency Effects (LFE).
[…]
Representation of our room setup.
@@
The goal of surround sound spatial audio generation is to generate ambisonic channels $\phi_{N}(t)$ from different non-spatial audios and positions $(x,y,z,\mathbf{A}^{T\times F})$. We utilize the pyroomacoustics library (Scheibler et al., 2018) to simulate sound propagation between sources and receivers. While the library implements a ShoeBox room model, our implementation focuses solely on direct sound propagation without computing room impulse responses (RIRs) or reflections, making the room geometry irrelevant to the final output. We place the listener centrally at the ground level, serving as a reference point for audio projection and perception. For the source and microphone placements, we adhere to the ITU-R BS 775 configuration (Series, 2010), which is essential for aligning our simulations with the standard 5.1 Surround Sound setups, including designated positions for the Front-Left (FL), Front-Right (FR), Center (C), Surround-Left (SL), Surround-Right (SR) and Low-Frequency Effects (LFE) channels.
Our implementation focuses on direct sound propagation, where the signal at each receiver position is determined by the geometric relationship between sources and receivers. For a source at position $\mathbf{s}$ and a receiver at position $\mathbf{r}$, the received signal is expressed as:
(4) $\phi_{ch}(t)=\sum_{i}\alpha_{i}\mathbf{A}_{i}^{T\times F}(t-\tau_{i}),$
where $\tau_{i}=\frac{\|\mathbf{r}-\mathbf{s}_{i}\|}{c}$ represents the propagation delay from source $i$ to the receiver, $c$ is the speed of sound, and $\alpha_{i}$ represents the attenuation due to distance. This model effectively captures the spatial distribution of sound sources without considering environmental reflections, making it suitable for both indoor and outdoor scenes.
@@
The synthesized output incorporates the interference patterns between multiple audio sources, which is a significant factor that differentiates our approach from scenarios where fewer sources are used and then combined. The resulting sound field exhibits unique characteristics that arise from the simultaneous interaction of all sources, creating spatial audio properties that are intrinsically linked to the arrangement $(x,y,z)$ and the number of sources in the complete setup.
5. Experiments and Results
@@
To the best of our knowledge, no previous techniques address the problem of spatial sound generation from visual content, thus we propose a novel quantitative evaluation technique based on marginal scene guidance as well as perform other qualitative evaluations (Section 5.1) alongside human evaluation (Section 5.2). We demonstrate qualitative examples in the supplementary.
5.1. Quantitative Evaluation
Table 1. Quantitative Evaluation. Scores are computed between each method and its AViTAR (Chen et al., 2022a)-enhanced version. Lower MFCC-DTW and higher ZCR, Chroma, and Spectral values denote greater similarity.
(+ AViTAR) DTW $\downarrow$ ZCR $\uparrow$ Chroma $\uparrow$ Spect $\uparrow$
CoDi (Tang et al., 2024) 0.0008 0.80 0.70 0.85
SEE-2-SOUND 0.00003 0.95 0.77 0.95
Table 2. Image–Audio CLIP Similarity.
Method CLIP $\uparrow$
Wav2CLIP (Wu et al., 2022) 0.0247
Im2Wav (Sheffer and Adi, 2023) 0.0108
SEE-2-SOUND 0.0259
@@ 5.1. Quantitative Evaluation
We present a new metric to conduct a quantitative evaluation of the scene guidance of generations produced by our model using various audio similarity metrics. We first sample images from Laion400M (Schuhmann et al., 2022) and the web. We then generate audio from these images using our method as well as CoDi (Tang et al., 2024) which serves as a baseline. We then use the audio outputs and images to generate modified scene-guided audio generated using AViTAR (Chen et al., 2022a). The use of an unconventional model-based evaluation stems from the lack of existing evaluations or any strong baselines for image-to-spatial audio. We report the average audio similarity scores between multiple such generated audio in Table 1. We particularly note that SEE-2-SOUND consistently generates spatial audio that matches the scene’s content often causing AViTAR (Chen et al., 2022a) to only slightly altering the generated audios. We show that SEE-2-SOUND performs significantly better in handling sound directiosn than other baselines.
We compute CLIP-score (Wu et al., 2022) to quantify the semantic similarty of the generated audio with the input image in Table 2. We sample 1000 images from the COCO (Lin et al., 2014) test set. While the CLIP score only measures the closeness of mono-audio with the image and our method spatial audio, we still demonstrate the CLIP score for SEE-2-SOUND. SEE-2-SOUND is comparable to other previous state-of-the-art methods on CLIP score.
5.2. Human Evaluation
Table 3. Human Evaluation. $\chi^{2}$-tests with $p<0.05$ indicate significant deviation from a uniform response distribution (lower is better).
Metric $p$-value $\downarrow$
Spatial-Audio Quality
@@ 5.2. Human Evaluation
   Overall localisation $4.523\times 10^{-4}$
   Direction detection $2.880\times 10^{-7}$
   Distance estimation $3.103\times 10^{-2}$
@@
Due to the absence of traditional metrics, we also use human evaluation to evaluate our approach. To promote future research in this field, we have open-sourced the evaluation protocol for human evolution. We conduct human evaluations by surveying $43$ human evaluators, for which we use randomly selected images and videos from the internet that are non-iconic in nature. We use this qualitative evaluation technique specifically due to the absence of any baseline methods for spatial generation. Additionally, the task of spatial audio generation is challenging since there are multiple possible ground truth spatial audios for a given image or video. This makes it difficult for us to use any kind of similarity metric. More details on the tasks in our user study can be found in our supplementary. We evaluate the results for each of these individual subtasks and questions and conducted $\chi^{2}$ tests to evaluate significance at the $p<0.05$ level.
@@
Spatial Audio Quality.:
@@
Adversarial Identification.:
Evaluators are given a random mix of generated spatial audio and mono audio. They are tasked with discriminating whether the given audio is indeed spatial audio, their accuracy is measured.
Audio Quality.:
@@
We summarize our results in Table 3. We particularly note that human evlauation users indicated a statistically significant preference towards SEE-2-SOUND in terms of spatial audio quality, directions of sound, and audio quality.
6. Discussion
@@
Lastly, inherent in the capabilities of generative models such as SEE-2-SOUND, which excel in producing synchronized and realistic outputs, is a significant concern regarding the potential generation and dissemination of deepfakes. The exploitation of this technology by malicious entities poses a notable risk even though we do not generate speech, as they could fabricate highly convincing counterfeit audio clips. We condemn any such application and hope to be able to work on safeguards for such approaches in the future.
7. Conclusion
@@
While not directly tested in this study, our work has significant implications for accessibility applications. The ability to generate spatial audio from visual content could potentially enhance the experience of visual media for individuals with visual impairments. Furthermore, this technology may have applications in medical settings, such as in neurology for patients recovering from stroke, and in broader accessibility and rehabilitation contexts. Future research could explore these potential applications to further understand and leverage the benefits of spatial audio generation in improving accessibility and quality of life for diverse user groups.
@@ Acknowledgements
The authors thank Bhavya Bhatt of the University of Toronto for his help on surveying the existing literature. This research was enabled in part by the support provided by the Digital Research Alliance of Canada  [^1] https://alliancecan.ca/. This research was supported in part with Cloud TPUs from Google’s TPU Research Cloud (TRC) [^2] https://sites.research.google/trc/about/. The resources used to prepare this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute  [^3] https://vectorinstitute.ai/partnerships/current-partners/.
@@ Appendix A Details of our Method
We provide details of our approach with which we demonstrate all of our experiments and results. Our experiments ran on a single A100-80 GPU. Running our approach on an image takes $\sim 12-15$ seconds (non-initial runs; initial runs can take up to $\sim 3$ minutes) for images and $\sim 5$ minutes on videos that are less than $20$ seconds like the ones we experiment with. Our method can be easily swapped for other off-the-shelf models, however we limit our experiments to these settings. Our human evaluation experiments were also performed without any text prompts to solely determine our framework’s performance.
@@
For source estimation we use a Segment Anything (Kirillov et al., 2023) model. We use the ViT-H/16 (Dosovitskiy et al., 2021) with $14\times 14$ windowed attention as our image encoder. Following standard approaches we rescale any inputs to $1024\times 1024$ by rescaling and padding. For the decoder, we use the same design and weights as used by Segment Anything (Kirillov et al., 2023). For monocular depth estimation, we use a Depth Anything (Yang et al., 2024) model. We use ViT-L/14 (Dosovitskiy et al., 2021) as our image encoder.
@@
For audio generation we use a CoDi (Tang et al., 2024) model and use the Video LDM, Audio LDM, and Text LDM components. For inference, we sample the model through $500$ diffusion steps.
@@
We use a sampling rate of $16000$ Hz for audio processing. For simulating the acoustic environment, we utilize a “ShoeBox” room model from Pyroomacoustics (Scheibler et al., 2018). The dimensions of the room are dynamically calculated based on the size of the image. Specifically, the dimensions width $w=$ image_width, height $h=$ image_height, and length $l=0.5\cdot$ image_width. In our virtual room, the microphones are placed at the following coordinates for each channel,
@@
Appendix B User Study Protocol
@@
B.1. Consent Form
Your decision to participate is voluntary, and you are free to withdraw at any time without any consequences. You are not required to answer any questions that you do not feel comfortable answering.
@@ B.2. Part 1: Familiarization
B.3. Part 2: Rating Spatial Audio with Images
@@
B.4. Part 3: Direction and Distance Identification
@@
B.5. Part 4: Real vs. Fake Identification
You will be shown 2 clips, one of them has spatial audio generated by and another one has mono audio of the same clips. You will be asked to identify which of them is ”real” spatial audio and ”fake” spatial audio.
@@
B.6. Part 5: Matching Images to Spatial Audio